# Object-centric proto-symbolic behavioural reasoning from pixels

### v1

@@ 4.2.3 Logical composition emergence
Figure 7: Logical composition. The OBR agent preference reasoning module is trained with two distinct rules and then tested in execution if the composed rule emerges. Task accuracy over 100 instances shows that OBR is combining both rules in the test phase.
@@ 5 Discussion
We described a brain-inspired deep learning architecture that allows grounded conditional abstract reasoning and behaviour by leveraging object-centric learned representation as the units of “mental” manipulation. Both perceptual and proto-symbolic goals were learnt in an unsupervised fashion from visual information (pixels). Results show that the system can learn and resolve conditional behaviours similar to 1st order logic: simple $A\rightarrow B$, complex $(A\rightarrow B\wedge C)\wedge(\neg A\rightarrow D\wedge E)$, XOR operation $(A\lor B\rightarrow C),A\wedge B\rightarrow D)$ and logical composition $(A\rightarrow B,A\rightarrow C\vdash A\rightarrow(B\wedge C)$. We analysed the model generalization to different number of objects (both in 2D and 3D), objects properties (e.g., color, shapes) and, critically, adaptation to changes in the environment (e.g., object substitution) thanks to the iterative inference and the preference network. OBR shows a promising direction in objects-based behavioral reasoning from pixels, mainly placing the foundations of grounded object-centric representations as a key inductive bias during learning, perception, resoning and control.
@@ 5.1 Grounded protosymbols and preferences in the latent space
While the agent behaviours from this work are much less complex from by current SOTA LLMs/VLM approaches such as RT architectures [48], researchers in the field are still struggling on how to perform goal-conditioned behavioural reasoning [49]. One of the key features of OBR is that it learns grounded representations and manipulate them in the latent manifold as well as it generates the agent preferences (goals) in the object-centric latent representation. This tights agent intentions to its own learning during interaction and constrained by its own learnt world model, bringing benefits but also restrictions as it forces the system to do embodied reasoning: the agents goals are tight to the world and the actions that it can perform in it.
5.2 Limitations and future work
Following the brain inspiration, we made some important decisions on the OBR architecture that derived into shortcomings. First, we employed an iterative inference procedure in order to benefit from postdictive inference (“smoothing”) in a sliding window approach. However, experimental results showed that the reasoner was highly dependent on the accuracy of the perceptual module. Thus, replacing the perceptual module by amortized methods could improve segmentation performance and reduce training time. Secondly, in the experiments, the effect of the real actions in the latent representation is assumed to be linear. While this could fit a hierarchical interpretation of cognition, it prevents to properly manipulate complex non-linear dynamics. While OBR can recover from collisions, it cannot plan ahead with them. Further work should investigate complex collisions, for instance, by incorporating a brain-inspired version of interaction networks. Furthermore, unsupervised learning of protosymbolic rules is dependant on the object learnt features. This implies that some combination of causal learnt behaviours are not properly resolved in execution. A holistic engineering solution for these limitations may be to incorporate a higher-layer that uses OBR as the interface between symbolic (e.g., graph-based, language) and subsymbolic representations. This will also expand the expressivity of the reasoning, which is currently limited, when comparing to LLMs reasoners. Finally, and very relevant to operationalize this approach for robotics applications the agent needs to include the body restrictions of performing an action in the world [50, 29]. This can be done by changing the action field mapping by a robotic arm controller.
@@ 6.1 Network architecture
OBR’s percpetual inference network consists of two separate Iterative Amortized Inference (IAI) modules, each of which in turn contains a refinement and a decoder module. The first IAI module concerns the inference of the state beliefs $q(\{{\bm{\mathbf{s}}}^{\dagger(k)}\})$ – we term this the perceptual inference module. The second IAI module infers the object action beliefs $q(\{{\bm{\mathbf{a}}}^{(k)}\})$, and we refer to this as the action inference module. In addition, OBR comprises a preference network that maps the current latent state beliefs to a predicted preference distribution over future object states. Code for the OBR architecture can be found at github.com/neuro-ai-robotics/OBR.

### v2

@@ 4.2.3 Logical composition emergence
Figure 7: Logical composition. The OBR agent preference reasoning module is trained with two distinct rules and then tested in execution if the composed rule emerges. Task accuracy over 512 instances shows that OBR is combining both rules in the test phase.
@@ 5 Discussion
We described a brain-inspired deep learning architecture that allows grounded conditional abstract reasoning and behaviour by leveraging object-centric learned representations as the units of “mental” manipulation. Both perceptual and proto-symbolic goals were learnt in an unsupervised fashion from visual information (pixels). Results show that the system can learn and resolve conditional behaviours similar to 1st order logic: simple $A\rightarrow B$, complex $(A\rightarrow B\wedge C)\wedge(\neg A\rightarrow D\wedge E)$, XOR operation $(A\lor B\rightarrow C),A\wedge B\rightarrow D)$ and logical composition $(A\rightarrow B,A\rightarrow C\vdash A\rightarrow(B\wedge C)$. We analysed the model generalization to different number of objects (both in 2D and 3D), objects properties (e.g., color, shapes) and, critically, adaptation to changes in the environment (e.g., object substitution) thanks to the iterative inference and the preference network. OBR shows a promising direction in objects-based behavioral reasoning from pixels, mainly placing the foundations of grounded object-centric representations as a key inductive bias during learning, perception, reasoning and control.
5.1 Object-centric representations as symbols
We showed that connectionist representation learning with inductive-bias, in particular object-centric representations, can work as a bridge between symbols and behaviour. Because objects are entities that live in both the real world and the internal representation of the agent, we can learn trajectories in the belief space that produce first-order-logic-like behaviours, where variables are substituted by object representations. Logical operators are encoded by learning the latent desired states (or preferences) of the agent and how to traverse the latent manifold to go from the current state to the desired state using the generative model.
@@ 5.2 Grounded protosymbols and preferences in the latent space
A key feature of OBR is that it learns grounded representations and manipulates them in the latent manifold, as well as generating the agent preferences (goals) in the object-centric latent representation. Agents’ internal goals are tied to the world and the actions that it can perform in it. Hence, bringing benefits but also restrictions as it forces the system to perform embodied reasoning. Learned protosymbolic rules depend on the potential learnable object features and how the agent can act on them—See latent space traversals in Appendix 8. For instance, the agent cannot learn to change the color of an object if all objects are in grey-scale or there is not a policy that can cause the desired effect. As a consequence, conditional behaviours that cannot be interpreted in the objects features space cannot be properly resolved in execution. On the good side, goal-conditioned behaviours are always grounded and therefore the gap between reasoning and control vanishes.
5.3 Shortcomings and future work
Following the brain inspiration, we employed an iterative inference procedure in order to benefit from postdictive inference (“smoothing”) in a sliding window approach. The resulting architecture is limited to the accuracy of the perception module. Thus, while this work focused on the proof-of-principle of the advantage of object-centric representations in behavioural reasoning, more powerful and sophisticated architectures (e.g. vision transformers) can be used to improve perceptual inference and scale to naturalistic images. Second, in the experiments, the effect of the real actions in the latent representation is assumed to be linear. While this could fit a hierarchical interpretation of cognition, it prevents to properly manipulate complex non-linear dynamics. While OBR can recover from collisions, it cannot plan ahead with them. Further work should investigate complex collisions, for instance, by incorporating other inductive biases, such as interaction networks [16] or sparsely changing latent representations to encode relevant events [48]. Third, the achieved agent behaviours are much less complex than by current SOTA LLMs/VLM approaches such as RT architectures [49]. OBR provides, however, goal-conditioned behavioural reasoning, which is a major challenge in the field [50]. Increased expressivity of the reasoning should be investigated, for instance, by using OBR as the interface between symbolic (e.g., graph-based, language) and subsymbolic representations. Finally, and very relevant to operationalize this approach for robotics applications the agent needs to include the body restrictions of performing an action in the world [51, 29]. This can be done by changing the action field mapping by a robotic arm controller.
@@ 6.1 Network architecture
OBR’s perceptual inference network consists of two separate Iterative Amortized Inference (IAI) modules, each of which in turn contains a refinement and a decoder module. The first IAI module concerns the inference of the state beliefs $q(\{{\bm{\mathbf{s}}}^{\dagger(k)}\})$ – we term this the perceptual inference module. The second IAI module infers the object action beliefs $q(\{{\bm{\mathbf{a}}}^{(k)}\})$, and we refer to this as the action inference module. In addition, OBR comprises a preference network that maps the current latent state beliefs to a predicted preference distribution over future object states. Code for the OBR architecture can be found at github.com/neuro-ai-robotics/OBR.